OpenStudy (anonymous):

SHOW ME HOW TO GET THE DERIVATIVE OF f(x) = 5x/(1+x^2). SHOW ME HOW TO DO IT. DON'T GIVE ME A HINT ASK WHAT I GET AND THEN SAY YOU DON'T KNOW HOW TO DO THAT, because then you're just BEING RETARDED and NO HELP WHATSOEVER THANK YOU

OpenStudy (rogue):

Use the quotient rule.\[\frac {d}{dx} \left[ \frac {u}{v} \right] = \frac {vu' - uv'}{u^2}\]\[\frac {df}{dx} = \frac {5(1 + x^2) - (5x)(2x)}{(1 + x^2)^2}= \frac {-5(x^2 - 1)}{(1 + x^2)^2}\]

OpenStudy (rogue):

Does that help?

OpenStudy (rogue):

Sorry, I meant \[\frac {d}{dx} \left[ \frac {u}{v} \right] = \frac {vu' - uv'}{v^2}\]

OpenStudy (rogue):

The denominator is supposed to be v^2 not u^2.

OpenStudy (anonymous):

YES it did perfectly thank you